# From #Dr00gtiktok to #harmreduction: Exploring Substance Use Hashtags on TikTok

### v1

@@ Methods
Network Analysis: For both the overall network and each community, we calculated standard network centrality measures including betweenness (Freeman 1977), closeness (Sabidussi 1966), degree (Freeman 1978), and eigenvector centralities (Bonacich 1972). These measures help identify influential hashtags and understand the structure of discourse within and across themes. Additionally, we computed network-level metrics such as average degree, clustering coefficient, and average shortest path length to characterize the overall connectivity patterns and community structure of substance use-related discourse on TikTok.
Characterizing Video Content
Codebook Development: To establish the codebook for coding substance-related TikTok videos, two researchers independently open-coded 160 randomly selected videos. Through this initial coding process, researchers identified nine distinct video topics and coded for substance presence and usage. The physical presence of drugs and related substances was of interest to us in determining which communities and topics were more engaged in depicting substances. The intercoder reliability was strong for coding substance visibility ($\kappa=0.87$) but only fair for video topics ($\kappa=0.35$). After discussing initial disagreements, the topics were refined from nine to seven distinct categories. The coders then had an iterative process for discussing the codes with all researchers on the team to identify coherent topics of videos within the TikTok dataset, resolve any disagreements, and finalize the codebook. The final codebook captured two key dimensions: video topic and substance presence, as shown in Table 3.
Video Selection: We started by selecting all videos that contained at least one of the 2,333 core hashtags, resulting in a dataset of 39,509 videos. To refine this dataset for detailed coding, we identified the top five hashtags per community based on their degree centrality.For each of the 80 selected hashtags, we then filtered videos using a stepwise approach:
  1. 1.
    50 initial videos were acquired from our dataset for each of the 80 hashtags. Ten of each were acquired from each quintile of viewership for the hashtag, achieving representation regardless of virality.
  2. 2.
    Each video was evaluated against three key criteria:
@@
    • Availability: The video is accessible and not moderated or removed by TikTok.
    • Language: The video features English as the primary spoken or written language.
@@
    • Substance Relevance: The video clearly depicts substance-related content, whether through visual cues, discussion, or behaviors.
We continued this process until we identified five videos per hashtag, one per each quintile of engagement, that met all three criteria. In total 977 links were coded, with 852 videos available (87.2%), of those 695 videos were in English (81.5%), and of those 351 videos were substance-related (50.5%). We conducted an inductive content analysis (Elo and Kyngäs 2008) on these 351 selected videos based on the codebook to ensure our final dataset was representative of the substance-use discourse across different communities (See Table 2).
Findings
@@
Table 2: Video statistics of the dataset filtered for RQ1 and the sub-set used for coding in RQ2.
RQ1: Network Structure and Communities
@@
Analysis of hashtags network showed complex, interconnected structures with distinct communities. The overall network contained 2,333 nodes (hashtags) and 46,990 edges (Table C in Appendix), with an average degree of 40.28 and a weighted degree of 513.84 as listed in Table C. These metrics indicate substantial interconnection - each hashtag co-occurs with approximately 40 other hashtags, and when hashtags do co-occur, they appear together frequently (over 500 times on average). The network exhibits small-world properties characterized by a high average clustering coefficient (0.52) coupled with a short average path length (2.20) (Watts and Strogatz 1998). This structure suggests that while substance-related discussions form tight communities, information can still flow efficiently across the network through just two intermediate hashtags on average.
Network centrality analyses revealed two distinct categories of influential nodes. The first category consists of substance-specific hashtags like “#addiction”, “#recovery”, “#sober”, “#wedorecover”, and “#harmreduction”, which scored highly across all centrality measures (betweenness, closeness, degree, and eigenvector). These hashtags function as critical bridges between different substance-related discussions, facilitating information flow between harm reduction, recovery, and other communities. The second category includes general-purpose hashtags like “#fyp” and “#mentalhealth”, which exhibit high centrality but serve primarily to connect substance-related discourse to broader TikTok conversations. This integration with platform-wide hashtags means substance-related content can potentially reach users not explicitly seeking such content through TikTok’s recommendation system.
Our community detection analysis identified 16 different themes (excluding Misc, see Table 1), each revealing distinct patterns in how substance-related content is organized on TikTok (Figure 6 in Appendix). The largest community in this network represents Awareness and Advocacy, containing approximately 12% of all hashtags (280 nodes) and featuring highly central nodes focused on harm reduction (e.g., “#harmreduction”, “#harmreductionworks”), recovery and sobriety (e.g., “#wedorecover”, “#sober”), and public health awareness (e.g., #opioidawareness, #opioidcrisis). The Emotions and Feelings cluster, while smaller with 180 nodes, shows high internal connectivity (clustering coefficient of 0.61) and frequently bridges to other communities through hashtags like “#mentalhealth” and “#anxiety,” suggesting the important role of emotional support in substance-related discussions. The network also revealed unique patterns in how substance-specific discussions are organized. Communities such as Cannabis (142 nodes) and Tobacco/Nicotine (138 nodes), exhibit strong internal clustering but limited external connections. This isolation suggests the discussions stay within their topical boundaries. The Commonly-Misused Substances cluster, though relatively small (120 nodes), contains hashtags often straddling the line between awareness and potential promotion of substance use (e.g., ”#studydrugs”, “#pingertok”, “#smartdrugs”). The Platform-specific cluster (165 nodes) showed distinct characteristics (e.g., “#fyp”, “#viral”), with the highest average betweenness centrality, indicating its crucial role in amplifying substance-related content to broader audiences.
@@
Figure 3: Engagement metrics across select popular hashtags in the coded set. Note, #harmreductionsaveslives has been shorted to HSL.
RQ2: Topics and Characteristics of Videos
Analysis of 351 videos based on our codebook (Table 3) showed that the most prevalent type of content in the videos focused on advocacy and recovery support (33.9%, 119 videos), where creators shared personal recovery journeys, celebrated sobriety milestones, or provided support to others struggling with substance use. For example, one video shows a creator celebrating “100 days sober milestone and encouraging others to seek help.” Satirical or humorous content formed the second largest category (28.2%, 99 videos), with videos often using comedy to address serious topics around substance use. The content is typically presented in the form of acted-out skits or parodies designed to gain engagement. For instance, one video presented a “comedic skit about stereotypical behavior while intoxicated.”
Informational content, typically from claimed medical or professional perspectives, comprised 13.1% (n=46) videos. These videos often featured “doctors explaining the effects of new medications on brain chemistry” or provided educational content about substance effects and risks. A smaller portion focused purely on social media engagement (3.9%, n=14 videos), with creators recording substance-related content primarily to gain followers or views. A sizable portion of the content is in the promotional category, appearing in 7% (n=24) videos. The majority of them portray a demonstration of a vaping device, Juul or lighter, and a few of them promote nootropics or study drugs. Notably, while discussions of substances were common, direct depiction of substances was less frequent - appearing in 25.9% (n=91) of videos, with only 23 of those videos showing active substance use. This suggests a tendency toward discussing rather than displaying substance use.
@@
Table 3: Codebook for qualitative analysis of substance-related TikTok videos (N=351)
Figure 4: Sub-networks of Awareness & Advocacy (Pink) and Health Conditions (Blue) to visualize strong overlap between hashtag usage.
@@
Alignment Between Hashtags and Video Content We also examined whether videos’ actual content aligns with their associated hashtag and communities. While each video was coded for its primary topic, we find that videos typically contain hashtags from multiple communities, revealing complex patterns in how substance-related content is tagged and disseminated.
@@
Advocacy, Recovery, and Support Content Videos coded primarily as advocacy and recovery content (119 videos, 33.9%), the largest thematic group in our manual coding, showed strong alignment with two major hashtag communities (see Table 4):
@@
  • Awareness & Advocacy (99 videos, 57.2%): Recovery-focused videos frequently used hashtags like “#harmreduction” and “#wedorecover”. Notably, while recovery-focused hashtags were more commonly used overall, content tagged with harm reduction messaging (particularly “#harmreductionsaveslives”) generated substantially higher user engagement across all metrics - views, likes, comments, and shares (Figure 3).
@@
  • Health Conditions (100 videos, 56.8%): Videos discussing recovery often included medical or health-focused tags like “#addiction”, “#chronicpain”, and “#pain” (Figure 4)
@@
Content of videos seen in this group are typically centered around active or past efforts to recover from substance abuse, discussed in context of cautionary tales, or implicitly imploring the viewer to regard substance abuse with regard and care due to potential harms.
@@
Entertainment and Platform Engagement Entertainment and humor-focused content (99 videos, 28.2%) as well as videos that documented using a substances for engagement (14 videos, 4%) similarly bridged multiple communities. These videos predominantly contained hashtags from the communities:
@@
  • Platform (82 videos, 40.6% of all coded videos): Entertainment videos heavily utilized platform-specific tags like “#fyp” and “#pingtok”
@@
  • Humor (60 videos, 78.9%): Videos using “#addictionhumor” and similar tags often presented substance use through comedy or satire
@@
  • Substance Effects (35 videos, 53.8%): Videos that were satirical or relatable often documented an effect of substance use (e.g., what it’s like being “#high” or “#stoned”). The subjects in these videos may not necessarily be intoxicated but could be enacting or portraying the effects.
Content in this group still discussed topics of substance use and misuse, however the seeming salient topic was humor. A large proportion of videos in this group were meme-based videos, where the original content was either not drug related but brought into drug discussion with filters, captions, or additional context, or was drug related but leaned towards humor and not object lessons.
@@
Figure 5: Distribution of video topics across three levels of substance visibility (no substances shown, substances shown but not used, and active substance use) in TikTok videos.
@@
Educational and Informational Content Videos coded as informational (46 videos, 13.1%) showed connections to multiple professional and health-focused communities:
  • Health Conditions (20 videos, 11.4%): Medical information and health impacts
  • Awareness & Advocacy (29 videos, 16.8%): Harm reduction education
  • Other Substances (16 videos, 24.6%) :
@@
  • Occupation (10 videos, 19.2%): Content from healthcare workers and other professions
Content in this group was thematically clinical or information focused. Humor was used sparingly if at all. The vast majority of videos in this cluster used detailed captions to go over specific facts regarding drug use or were more long-form and had a health professional or recovery advocate speak at length on the topic. Recovery organizations were seen multiple times in the group, with high production value videos.
@@
While advocacy, entertainment, and educational content showed strong alignment with specific hashtag communities, certain clusters served as contextual bridges across all content types. Hashtags from the Commonly Misused Substances (111 videos, 31.6%), Identity & Community (n = 95 videos, 27.1 %), and Occupation (53 videos, 15.1 %) clusters appeared across all seven content categories, suggesting their role in providing context rather than signaling specific content types. For instance, videos tagged with “#opiates” spanned multiple content categories: healthcare workers explaining naloxone administration (Recovery/Advocacy), recovery nurses humorously discussing their reliance on caffeine (Satirical), and medical professionals explaining medication effects (Informational). Similarly, occupation-related hashtags like “#nurse” or “#doctor” and identity-based tags like “#disabled” or “#veteran” appeared across all content. Consumption Method hashtags like “#vape” or “#smoke” (101 videos, 28.8% of videos) served a more specific bridging function, describing substance use behaviors across content types.
RQ3: Effects of Substance Visibility

### v2

@@ 3.2.2. Network Analysis:
For both the overall network and each community, we calculated standard network centrality measures including betweenness (Freeman, 1977), closeness (Sabidussi, 1966), degree (Freeman, 1978), and eigenvector centralities (Bonacich, 1972). These widely used measures capture complementary aspects of structural prominence, including reachability, connectivity, and association with highly connected hashtags and help identify influential hashtags in terms of their structural roles within the discourse network. Additionally, we computed network-level metrics such as average degree, clustering coefficient, and average shortest path length to characterize overall connectivity, cohesion, and small-world properties of substance use–related discourse on TikTok, both at the global network level and within individual communities.
3.3. Qualitative Analysis of Video Content (RQ2)
Building on the hashtag communities identified in RQ1, we analyzed video content associated with the 2,333 core hashtags organized into 16 thematic clusters (excluding the Misc cluster). While network analysis reveals how substance-related discourse is structurally organized, it cannot capture the nature of content within these communities (e.g., whether videos promote use, support recovery, or serve other functions). To address RQ2, we qualitatively examined how substance-related discourse manifests within and across these communities along two dimensions: video topic and substance presence.
3.3.1. Codebook Development
We developed a qualitative codebook using an iterative open-coding process designed to capture salient characteristics of substance-related video content (braun2006using). Two members of the research team independently coded an initial set of 160 randomly selected videos drawn from the broader 39,509-video dataset. Through this exploratory open-coding process, the coders identified an initial set of nine video topic categories reflecting recurring themes in the content. In parallel, we coded for substance presence, capturing whether substances were visually depicted or actively used within videos. This dimension was included to assess how different communities and topics varied in their explicit depiction of substances. Following this initial round, the research team met to discuss discrepancies, refine category definitions, and consolidate overlapping themes. Through another round of independent coding, reliability was assessed on a 10% (n = 40) of the coded videos, yielding moderate agreement for video topic ($\kappa=0.51$) and substance shown ($\kappa=0.58$). This process resulted in a reduced set of seven video topic categories. Remaining disagreements were resolved through deliberative discussion among the full research team until consensus was reached. The 160 development videos were excluded from the final analytic sample to maintain independence between codebook development and application. The final codebook captured two primary dimensions: (1) video topic and (2) substance presence. Definitions and examples for each category are provided in Table 3.
3.3.2. Video Sampling
We constructed the qualitative sample through a multi-stage, stratified selection process designed to ensure coverage across communities and engagement levels. We started with the full set of 39,509 videos containing at least one of the 2,333 core hashtags. To ensure representation across the 16 hashtag communities, we first selected the top five hashtags per community based on degree centrality, yielding a total of 80 focal hashtags. Degree centrality was used to prioritize hashtags that were structurally prominent within each community’s discourse network. For each focal hashtag, we stratified associated videos into five engagement quintiles based on view count, calculated within each hashtag’s distribution.
Our target was to code at least one video per engagement quintile per hashtag (5 videos × 80 hashtags = 400 minimum). To account for anticipated exclusions due to content removal, non-English language, or lack of substance relevance, we initially sampled up to 10 videos per quintile where available. However, many hashtags had fewer than 50 associated videos in the filtered dataset, resulting in incomplete coverage of some quintile-hashtag combinations. This process yielded 977 candidate videos for screening. Each candidate video was then evaluated against three inclusion criteria:
@@ 3.3.2. Video Sampling
  • Availability: The video was publicly accessible at the time of coding (i.e., not removed or restricted).
  • Language: The primary spoken or written language of the video was English.
@@
  • Substance Relevance: The video clearly involved substance-related content, indicated through visual cues, discussion, or behaviors.
Of the 977 candidate videos, 852 (87.2%) were publicly available at the time of analysis, 695 (81.5% of available videos) were in English, and 351 (50.5% of English videos) met the substance relevance criterion. These 351 videos constituted the final corpus for inductive qualitative content analysis (Elo and Kyngäs, 2008). The videos were divided between the two coders for independent coding. Coders consulted with the broader research team when encountering ambiguous cases to ensure consistent application of the codebook. We report descriptive frequencies for all coded dimensions. Descriptive statistics for the sampled videos are reported in Table 2.
4. Findings
@@
Table 2. Video statistics of the dataset filtered for RQ1 and the sub-set used for coding in RQ2.
4.1. RQ1: Network Structure and Communities
@@
Analysis of the hashtag network revealed a highly interconnected structure with 2,333 nodes (hashtags) and 46,990 edges (Table A.1 in Appendix), with an average degree of 40.28 and a weighted degree of 513.84. These metrics indicate substantial interconnection; each hashtag co-occurs with approximately 40 other hashtags, and when hashtags do co-occur, they appear together frequently (over 500 times on average). The network exhibits small-world properties (watts_collective_1998), characterized by a high average clustering coefficient (0.52) and a short average path length (2.20). This structure suggests that substance-related discussions form tightly knit clusters while remaining structurally connected through relatively short paths across the broader network.
Network centrality analysis identified recovery-focused hashtags as structurally central. The five hashtags with highest betweenness centrality: #addiction (0.129), #recovery (0.098), #sober (0.058), #wedorecover (0.050), and #harmreduction (0.049), all relate to recovery and harm reduction, positioning these terms as bridges connecting otherwise disparate thematic communities.
Community analysis identified 16 thematic clusters (excluding Misc, see Table 1), each reflecting distinct structural properties within substance-related discourse (Figure A.1 in Appendix). Communities varied substantially in internal cohesion. The largest community, Awareness and Advocacy, comprises approximately 12% of all hashtags (279 nodes) with high clustering (0.89), and includes hashtags related to harm reduction (e.g., #harmreduction, #harmreductionworks), recovery and sobriety (e.g., #wedorecover, #sober), and public health awareness (e.g., #opioidawareness, #opioidcrisis). Similarly, Health Conditions (277 nodes, clustering = 0.85) showed high internal cohesion around hashtags such as #addiction and #chronicpain.
Substance-specific communities exhibit varied structural patterns. Cannabis (51 nodes, clustering = 0.55) and Tobacco/Nicotine (31 nodes, clustering = 0.70) show high internal clustering, while Commonly-Misused Substances (152 nodes, clustering = 0.38) and Alcohol (43 nodes, clustering = 0.41) exhibit more moderate cohesion. The Commonly-Misused Substances community contains hashtags that often span awareness-oriented and substance-use-oriented content (e.g., #studydrugs (petersen_studydrugspersuasive_2021), #pingertok (Whelan et al., 2024), #smartdrugs (petersen_studydrugspersuasive_2021)).
In contrast, several communities function as cross-cutting connectors with low internal clustering. The Consumption Method community (41 nodes, clustering = 0.30), containing hashtags such as #smoke, #vape, and #injection, exhibit the highest average betweenness centrality among all communities, connecting discussions across substance types. The Platform-Specific community (183 nodes, clustering = 0.12), characterized by hashtags such as #fyp, #viral, exhibits minimal internal structure with only 3.2% of its edges occurring within the community, consistent with its function as a mechanism for algorithmic visibility rather than topical organization.
@@
Figure 3. Engagement metrics across select popular hashtags in the coded set. Note, #harmreductionsaveslives has been shorted to HSL.
4.2. RQ2: Video Topics and Substance Presence
Analysis of 351 videos using the qualitative codebook (Table 3) revealed that the most prevalent category of content focused on advocacy and recovery support (33.9%, $n=119$ videos). Videos in this category commonly featured creators sharing personal recovery journeys, celebrating sobriety milestones, or offering encouragement and support to others navigating substance use challenges. For example, one video documented a creator celebrating a “100 days sober” milestone while encouraging viewers to seek help.
Satirical and humorous content comprised the second largest category (28.2%, $n=99$ videos). These videos frequently used comedy or parody, often through skits or reenactments (e.g., “comedic skit about stereotypical behavior while intoxicated”), to engage with substance-related experiences. Informational content, often framed from claimed medical or professional perspectives (e.g., “doctors explaining the effects of new medications on brain chemistry”), accounted for 13.1% ($n=46$) of videos and included educational explanations of substance effects, risks, or treatment options. A smaller subset of videos focused primarily on platform engagement (3.9%, $n=14$ videos), where substance-related themes were used mainly to attract views or followers. Promotional content appeared in 7.0% ($n=24$) of videos, most commonly featuring demonstrations of vaping devices, Juul or lighter, as well as occasional promotion of nootropics or so-called study drugs. Across all content categories, direct depiction of substances was comparatively limited. Substances were visually present in 25.9% ($n=91$) of videos: 19.4% ($n=68$) showed substances without active use, while only 6.6% ($n=23$) depicted active substance use. This pattern suggests that substance-related discourse on TikTok more frequently involves discussion, narration, or symbolic reference rather than explicit visual display.
@@ 4.2. RQ2: Video Topics and Substance Presence
Table 3. Codebook for qualitative analysis of substance-related TikTok videos (N=351)
@@
We next examined how videos’ content aligned with their associated hashtag and communities. While each video was coded for its primary topic, videos typically contained hashtags from multiple communities, revealing complex patterns in how substance-related content is tagged and disseminated. The following analysis focuses on the three most prevalent content categories: Advocacy and recovery, Entertainment, and Educational content - which together comprise 75.2% of coded videos. Full cross-tabulations for all categories are provided in Table A.2 in the Appendix.
@@
Advocacy, Recovery, and Support Content: Videos coded primarily as advocacy and recovery content (33.9%, $n=119$ videos), the largest thematic group in our manual coding, showed strong alignment with two major hashtag communities (see Table A.2 in Appendix):
@@
  • Awareness & Advocacy (57.2%, $n=99$ videos): Recovery-focused videos frequently used hashtags like #harmreduction and #wedorecover. Notably, while recovery-focused hashtags were more commonly used overall, content tagged with harm reduction messaging (particularly #harmreductionsaveslives) generated substantially higher user engagement across all metrics - views, likes, comments, and shares (Figure 3).
@@
  • Health Conditions (56.8%, $n=100$ videos): Videos discussing recovery often included medical or health-focused tags like #addiction, #chronicpain, and #pain (Figure 4)
@@
Content in this group was typically centered on active or past efforts to recover from substance use, framed through cautionary narratives or expressions of care regarding the potential harms of substance misuse.
Figure 4. Sub-networks of Awareness & Advocacy (Pink) and Health Conditions (Blue) to visualize strong overlap between hashtag usage.
@@
Entertainment and Platform Engagement Entertainment and humor-focused content (28.2%, $n=99$ videos) as well as videos that documented using substances for engagement (4%, $n=14$ videos) also bridged multiple communities. These videos predominantly contained hashtags from the following communities:
@@
  • Platform (40.6%, $n=82$ videos): Entertainment videos heavily utilized platform-specific tags like #fyp and #pingtok
@@
  • Humor (78.9%, $n=60$ videos): Videos using #addictionhumor and similar tags often presented substance use through comedy or satire
@@
  • Substance Effects (53.8%, $n=35$ videos): Videos that were satirical or relatable often documented an effect of substance use (e.g., what it’s like being #high or #stoned). The subjects in these videos may not necessarily be intoxicated, but could be enacting or portraying the effects.
While substance use and misuse remained a topic in these videos, humor served as the dominant framing. A large portion of videos in this group were meme-based, in which non–drug-related content was reframed through captions, filters, or contextual cues, or drug-related content was presented primarily for comedic effect rather than instructional or cautionary purposes.
@@
Figure 5. Distribution of video topics across three levels of substance visibility (no substances shown, substances shown but not used, and active substance use) in TikTok videos.
@@
Educational and Informational Content Videos coded as informational (13.1%, $n=46$ videos) showed connections to multiple professional and health-focused communities:
  • Health Conditions (11.4%, $n=20$ videos): Medical information and health impacts
  • Awareness & Advocacy (16.8%, $n=29$ videos): Harm reduction education
  • Other Substances (24.6%, $n=16$ videos) :
@@
  • Occupation (19.2%, $n=10$ videos): Content from healthcare workers and other professions
Content in this group was predominantly clinical or informational in tone. Humor was used sparingly, if at all. Most videos relied on detailed captions or longer-form explanations, often delivered by health professionals, recovery advocates, or organizations, many of which exhibited higher production value.
@@
Cross-Cutting Communities. While advocacy, entertainment, and educational content showed strong alignment with specific hashtag communities, certain clusters served as contextual bridges across all content types. Hashtags from the Commonly Misused Substances (31.6%, $n=111$ videos), Identity & Community (27.1 %, $n=95$ videos), and Occupation (15.1 %, $n=53$ videos) clusters appeared across all seven content categories, suggesting their role in providing context rather than signaling specific content types. For instance, videos tagged with #opiates spanned multiple content categories: healthcare workers explaining naloxone administration (Recovery/Advocacy), recovery nurses humorously discussing their reliance on caffeine (Satirical), and medical professionals explaining medication effects (Informational). Similarly, occupation-related hashtags like #nurse or #doctor and identity-based tags like #disabled or #veteran appeared across all content. Consumption Method hashtags like #vape or #smoke (28.8%, $n=101$ videos) served a more specific bridging function, describing substance use behaviors across content types.
4.3. RQ3: Effects of Substance Visibility